OpenStudy (anonymous):

Can someone please explain in detail how to solve (3ab)(-5ab) using distributing? Like, in a lot of detail. It's my first time doing a math problem in a LONG time.

OpenStudy (anonymous):

sorry give me a moment and i ll explain u

OpenStudy (sburchette):

With this are multiplying two monomials. You can combine variables with like bases by using exponent rules. First you can multiply the coefficients 3 and -5 to get -15. So now we have \[-15(ab)(ab)\] To multiply the variables, you will use the exponent rule for multiplication which is \[(a^n)(a^m)=a ^{m+n}\] Meaning that you add the exponents if the bases are the same. So in the case of a*a you have \[a^1*a^1=a^{1+1}=a^2\] And with the b variables you have \[b^1*b^1=b^{1+1}=b^2\] So combining all of these parts we end up with \[-15a^2b^2\]